OpenStudy (anonymous):

assume that x and y are both differentiable functions of t and find the required values of dy/dt and dx/dt dx/dt when x=4 dx/dt=3

OpenStudy (anonymous):

$$\frac{dx}{dt} = 3 \\ dx = 3 dt \\ \int dx = \int 3 dt = 3t$$ Where do you go from here?

OpenStudy (anonymous):

I forgot to say that the function is the square root of x

OpenStudy (anonymous):

Which function?

zepdrix (zepdrix):

\[\Large x=x(t)\]\[\Large y=y(t)\] \[\Large y=\sqrt x\] I don't quite understand your question. Is this what you meant to write? `Find dy/dt when x=4 and dx/dt=3`

OpenStudy (anonymous):

Yes using the square root of x I got that dy/dt= 3/4

zepdrix (zepdrix):

Hmm yes that's what I'm coming with also. Good job! :)\[\Large \frac{dy}{dt}\quad=\quad \frac{1}{2\sqrt x}\frac{dx}{dt}\]Plugging in the values appears to give us 3/4!

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.